OpenStudy (anonymous):

If r(x) = 3x − 1 and t(x) = −4x − 1, find (r + t)(7). Do anyone know the answer to that?

OpenStudy (helder_edwin):

just use the definition \[ \large (r+t)(7)=r(7)+t(7) \]

OpenStudy (anonymous):

so its 49?

OpenStudy (anonymous):

ok thanks.

OpenStudy (helder_edwin):

let's see \[ \large (r+t)(7)=r(7)+t(7)=(3\cdot7-1)+[(-4)\cdot7-1]= \] \[ \large =(21-1)+(-28-1)=20+(-29)=-9 \]

OpenStudy (anonymous):

how about this one If f(x) = 1 over 3x − 2 and g(x) = −2x − 3, find (f • g)(6)

OpenStudy (anonymous):

Would it be -15?

OpenStudy (helder_edwin):

it is pretty much the same \[ \large (f\cdot g)(6)=f(6)\cdot g(6) \]

OpenStudy (helder_edwin):

\[ \large =\frac{1}{3\cdot6-2}\cdot(-2(6)-3)=\frac{1}{18-2}\cdot(-12-3) \] \[ \large =\frac{1}{16}\cdot(-15) \]

OpenStudy (anonymous):

thanks

OpenStudy (helder_edwin):

u r welcome. did u get it?

OpenStudy (anonymous):

kinda